# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ Methods
This concludes the description of the computational model. A brief overview over the parameters an their effects, several of which will be discussed in detail below, is given in SI Table 1 and Fig. S8. Fig. S15 provides a flow chart of the simulation algorithm.
@@ Results
Figure 2: Extent of tissue flow depends on initial cell-scale order and tension anisotropy. A, B Simulation snapshots of tissue patches with free boundaries starting from an (irregular) hexagonal cell array (A) and a disordered cell array (B) generated from a random Voronoi tessellation. The net deformation is significantly smaller in the latter case. In both cases the final configuration is disordered. Each snapshot shows the top half of the tension triangulation and the bottom half of the corresponding cell array. Note the initial tension triangulation is stretched along the $x$ direction and compressed along the $y$ direction as a result of the prescribed initial tension anisotropy. C Convergent extension (measured by the tissue aspect ratio $a(t)$) is slower and ceases at a smaller net deformation for a disordered initial condition (cyan line) compared to the initially ordered case (purple line). Shaded bands indicate standard deviation over $N=3$ simulation runs with $N_{\mathrm{cells}}\approx 10^{3}$ cells each. D Plotting the rate of tissue convergent extension against the T1 rate provides a measure for the efficiency of T1 transitions. For the initially ordered tissue, the efficiency of T1s is starts out near the theoretical optimum but drops to zero as the tissue becomes disordered. For the disordered initial condition, T1s are inefficient from the beginning. The theoretical bound (for cells retaining isotropic shape) is a factor of $\sqrt{3}$ elongation for one T1 per cell (i.e. each cell participates in two T1s) as illustrated in the inset cartoon. E Decreasing hexagon fraction (solid lines) and global tension anisotropy (dashed lines) indicate the decay of order in the cell arrays. Notably, in the initially ordered cell array, tension anisotropy transiently increases due to positive tension feedback (solid purple line). At late times both simulations converge to zero global tension anisotropy and identical coordination number statistics (inset). F As order is lost, the cell-shape index $s$ converges to that of a random Voronoi tessellation $s_{\mathrm{V}}$ (dashed black line). Note that the critical shape index for the rigidity transition in the “classical” vertex model [29], $s_{0}^{*}\approx 3.81$ (dot-dashed orange line), does not feature in our results. Lines show the median, shaded bands indicates quartiles.
@@
In the companion paper [17], we have shown that positive tension feedback can drive an active T1 transition in a cell quartet where the inner interface is initially under slightly higher tension than its neighbors. The simulations were performed in a quartet of identical cells, representing a perfectly regular lattice of cells. However, any real tissue will exhibit some degree of irregularity. Investigating the effect of this disorder is key to understand convergent extension on the tissue scale. To this end, we perform simulations of irregular cell arrays. All parameters are set to the same values as in the companion paper, where the time scale $\tau_{T}$ was calibrated to fit the tension and interface length dynamics of active T1s during Drosophila gastrulation. The local T1-dynamics of the tissue-scale simulations are quantified in Fig. S5, and match the experimental data quite well. We start with simulations of a freely suspended patch of tissue to investigate the role of initial tension anisotropy and order in the cell packing. Further below, we will present simulations that combine active with passive tissue regions, mimicking the Drosophila germ band and the adjacent amnioserosa tissue.
[…]
@@
The tissue flow is driven by self-organized cell rearrangements (active T1 transitions) whose rate rapidly increases, reaching a maximum, and then decreases to a lower, but non-zero, value (Fig. 2D). Large-scale tissue deformation stalls after approximately 2-fold convergent extension (as measured by the square root of the aspect ratio $a=\mathrm{width}/\mathrm{height}$, Fig. 2C) while cells continue rearranging. This suggests that T1s at this stage are no longer coherently oriented and therefore do not contribute to net tissue deformation. Indeed, the efficiency of T1s at driving convergent extension drops after the T1 rate has peaked (Fig. 2D).
@@
To test the role of initial order in the cell packing, we initialize the simulation from a random Voronoi tessellation but an identical tension anisotropy of $0.175$, which results in slower convergent-extension flow and arrest of flow at a smaller amount of total tissue-scale deformation (Figs. 2B and 2C and Movie 2). Notably, tension anisotropy rapidly vanishes without the transient increase observed in the simulation starting with a more ordered cell packing (Fig. 2E). While the early dynamics depends sensitively on the initial condition, we find rapid convergence toward a common disordered steady state.
A common measure for the degree of order in a polygonal tiling is the cell shape index ($s=P/\sqrt{A}$, with cell perimeter $P$ and area $A$), shown in Fig. 2F. This shape index is high when cell shapes are elongated or irregular and approaches the minimal value $s\approx 3.72$ for a regular hexagon. The shape index has a particular relevance in vertex models employing an area-perimeter based elastic energy [30, 7]. In these models, the target shape index is a control parameter that drives a solid to fluid transition [29]: a high target shape index ($s_{0}>s_{0}^{*}\approx 3.81$) is associated with tissue fluidity since it allows for cell rearrangements, while a low cell target shape index gives rise to a solid state. In both cases, the observed shape index is controlled by the target shape index. In contrast, in our simulations with actively driven T1s, disorder, and thus a high observed shape index, is the consequence of cell rearrangements, rather than their cause. Notably, we find more tissue flow when the observed shape index is initially low (Fig. 2F). The question of the solid vs fluid character of tissue will be addressed in more detail in the last part of the results section and in the discussion.
To systematically investigate the role of tension anisotropy and cell packing order on self-organized convergent extension, we generated initial tension triangulations from random hard disk packings at different packing fractions $\rho$ [31, 32]. Specifically, we sample the positions of the tension vertices $\mathbf{t}_{i}$ from a hard disk process, and then construct the corresponding Delaunay triangulation. At low packing fraction, the hard disk process generates highly irregular triangulations with broadly distributed coordination numbers (see Fig. 2B). In contrast, at sufficiently high packing fraction $\rho\gtrsim 0.72$, the disks adopt a crystalline hexagonal packing such that $p_{6}\approx 1$. The plot at the bottom of Figure 3A shows how the degree of order continuously varies with $\rho$ between these two extremes. To introduce a specified initial tension anisotropy, the triangulation is sheared with magnitude $s$ (displacing vertices by $\mathbf{t}_{i}\mapsto\mathrm{diag}(\sqrt{1+s},1/\sqrt{1+s})\,\mathbf{t}_{i}$).
Figure 3: Hexagon fraction and tension anisotropy predict tissue flow. A The total amount of convergent-extension (overall tissue aspect ratio change $a_{\mathrm{final}}/a(0)$) increases as a function of cell packing order (controlled by hard disk packing fraction $\rho$ used to generate the Voronoi seed points) and initial tension anisotropy (controlled by a strain $s$ applied to the tension triangulation). The hard disk packing fraction $\rho$ controls the fraction of hexagonal cells (bottom). The purple and green stars indicate the initial condition parameters used for Fig. 2A and Fig. 2B respectively. The red star indicates the approximate state of the Drosophila germ band at the onset of germ band extension (cf. Fig. S1B). B Hexagon fraction, $p_{6}$, and tension anisotropy $||\langle\sqrt{2}\tilde{Q}\rangle||$ decrease as cells rearrange and eventually converge to a fixed point where anisotropy vanishes and the hexagon fraction matches a random Voronoi tessellation ($\sim 0.4$). At this point, convergent-extension flow, as measured by the remaining change in tissue aspect ratio $a_{\mathrm{final}}/a(t)$, stalls (cf. Fig. 2C, E). The trajectories approximately collapse onto a plane, suggesting that a linear relationship can predict the convergent-extension in terms of $p_{6}$ and $||\langle\sqrt{2}\tilde{Q}\rangle||$; see Eq. (6). C Correlation between predicted and observed change in aspect ratio. Data points are from simulations with initial conditions, spanning the entire parameter diagram in (A), showing that the prediction works across many different initial conditions. Points with the same color indicate different time point from the same simulation.
The heatmap in Fig. 3A (top) shows the dependence of convergent extension on the initial configuration controlled by $\rho$ and $s$. The total extent of convergent extension, quantified by the net change in aspect ratio $\sqrt{a_{\mathrm{final}}/a_{\mathrm{initial}}}$ increases continuously with both the initial order and the magnitude of tension anisotropy. As we have seen above, both these quantities decrease as cells rearrange (cf. Fig. 2E). In fact, when plotting the remaining extent of convergent extension $\sqrt{a_{\mathrm{final}}/a(t)}$ against $p_{6}(t)$ and $||\langle\sqrt{2}\tilde{Q}(t)\rangle||$, we find trajectories that approximately lie in a common plane and converge to a fixed point at vanishing anisotropy $||\langle\sqrt{2}\tilde{Q}\rangle||\approx 0$ and $p_{6}\approx 0.4$ (Fig. 3B). Based on these results, we hypothesize an empirical law for feedback-driven convergent extension, based on the instantaneous hexagon fraction and tension anisotropy:
$\sqrt{\frac{a_{\mathrm{final}}}{a(t)}}\approx c_{Q}||\langle\sqrt{2}\tilde{Q}(t)\rangle||+c_{6}p_{6}(t)$ (6)
With the coefficients determined by a linear fit, Eq. (6) predicts the remaining extent of convergent extension with less than 10% mean absolute error (Fig. 3C). The empirical relation Eq. (6) is specific to the particular choice of microscopic tension dynamics employed in the simulations: a systematic search for a more general relation linking microscopic ($p_{6},||\langle\sqrt{2}\tilde{Q}(t)\rangle||$) and macroscopic ($a(t)$) quantities is beyond the scope of the present work. Note further that we have focused on degree of order in the topology (i.e. coordination numbers) of the triangulation, quantified by $p_{6}$. A different manner of introducing disorder is via random displacements of the triangulation vertices while keeping the topology fixed, which we study in (see Fig. S8). This form of disorder also decreases convergence-extension flow.
The model introduced above implements a specific form of positive tension feedback, Eq. (5), in such a way that total tension is conserved locally. We find qualitatively similar results for tension dynamics with different local conservation laws. Notably, positive feedback conserving areas of tension triangle instead of their perimeters produces slightly more convergent extension and exhibits an even stronger dependence of tissue flow on the initial degree of order (see SI Sec. 2). How tissue dynamics depends on the parameters $\tau_{\mathrm{balance}},\mu,\lambda,n$ is summarized in SI Table 1 and Fig. S8. We will return to the effect of specific parameters throughout the following sections.
Taken together, our model reproduces several salient features of germ-band extension in the Drosophila embryo. We find that convergent extension driven by positive tension feedback is self-limiting and naturally explains the transition from the fast to the slow phase of germ-band extension [17]. It also reproduces that the slowdown of convergent extension is concomitant with an increase in cell-scale disorder, approaching a maximally disordered state [33, 17]. Our model suggests that the transient fast phase of flow is facilitated by an initial near-hexagonal packing of the cells. In the Drosophila embryo, this packing results from physical interactions of the spindle apparatuses during the syncytial cell division cycles that precede cellularization [34, 35]. Quantification of hexagonal order in the experimental data are show in Fig. S1. We predict that disrupting the initial cell packing will cause GBE to become slower. The experimental data analysis in Ref. [17] demonstrates the presence of large-scale tension anisotropy before the onset of convergent extension, which in the model is required to orient tissue flow. More precisely, Ref. [17] as well as Figs. S1, S2, and S3 suggest that the Drosophila germ band corresponds, approximately, to a hard disk packing fraction $0.65$ and an initial anisotropy of $0.3$ on the convergent extension heat map Fig. 3A. This predicts an ca. $1.85$-fold extension. We attribute the additional elongation observed in the experimental data to external pulling forces [8] and ventral furrow invagination, which provides an external source of tension anisotropy via strain-rate feedback [14]. Both factors are not present in the simulations Fig. 3A. More broadly, our results suggest that convergence extension via positive tension feedback may generally be self-limiting, as also noted in a recent model by Sknepnek et al.[10].
@@
So far, we have focused on the role of anisotropy and initial topological order in the cell packing. In the Drosophila germ band, we additionally observed a more subtle form of geometric order – a particular pattern of alternating high and low tensions [17] – that precedes the onset of cell rearrangements. This cell-scale tension pattern is not apparent initially but arises dynamically in conjunction with an increase in overall tension anisotropy which suggests that this tension pattern arises from the positive tension feedback. As we show in the following, systematically quantifying the local configurations of cortical tensions provides a method to constrain physical models of cell rearrangements in epithelia.
@@
Figure 4: Triangle shape space characterises the local tension motifs that underlie cell rearrangements. A We distinguish two tension patterns, “cables” and “bridges”, by their elementary motifs of tension configuration at individual vertices. These configurations can be quantified in terms of the tension triangle shapes which we use to define a local tension configuration (LTC) parameter. B Shape decomposition of a triangle into a sequence of three transformations acting on an equilateral reference triangle. The angle $\psi$ and the shear factors $s_{1,2}$ determine the intrinsic shape of the triangle while the angle $\phi$ determines its orientation in space. C The intrinsic shape of triangles can be quantified by a complex order parameter $\tilde{s}\,e^{i\tilde{\psi}}$, where the LTC magnitude $\tilde{s}$ is proportional to the triangle’s anisotropy and the LTC phase $\tilde{\psi}$ distinguishes obtuse and acute triangles. The gray line indicates the Delaunay condition for a pair of identical triangles. Along this line, the circumcenters of the two triangles coincide, corresponding to a fourfold vertex in the Voronoi tessellation. This line therefore defines a threshold for T1 events in the case of vanishing isogonal strain. D Circumcircle construction of the Voronoi edge length (purple) from a pair of adjacent (tension) triangles (red). Circumcircles are indicated by gray dashed lines. The actual physical length $\ell$ is the sum of the Voronoi reference length $\ell_{\mathrm{ref}}$, and a contribution from isogonal strain $\Delta\ell_{\mathrm{iso}}$. In the illustration, $\Delta\ell_{\mathrm{iso}}/\ell_{0}$ is negative. E T1 threshold as a function of the isogonal strain $\Delta\ell_{\mathrm{iso}}/\ell_{0}$. Positive isogonal strain shifts the threshold to higher tension anisotropy.
The elementary motifs of a cell-scale tension pattern are the tension configurations at individual vertices. In force balance, the tension vectors at each vertex form a triangle. This allows one to characterize the tension configuration based on the shape of the tension triangle: acute triangles correspond to tension cables (adjacent high tension interfaces) while obtuse triangles correspond to an isolated high tension interface which we refer to as a tension “bridge”. The latter configuration is the elementary motif of an alternating pattern of high and low tensions as illustrated in Fig. 4A.
@@
To quantify the relative abundance of these motifs, we define a local tension configuration (LTC) parameter that measures how anisotropic and how acute vs obtuse the tension triangle is. We order the three tension vectors $\mathbf{T}_{\alpha},\alpha=1,2,3$ that form the tension triangle by increasing length, i.e. $T_{1}\leq T_{2}\leq T_{3}$. We then use the force-balance condition $\sum_{\alpha}\mathbf{T}_{\alpha}=0$ to define the reduced barycentric vectors, combined into a 2$\times$2 matrix
@@
where the normalization factor $\mathcal{N}$ ensures $||\mathfrak{T}||^{2}=\mathrm{Tr}[\mathfrak{T}\mathfrak{T}^{T}]=1$. This normalization fixes the arbitrary overall tension scale. Note that $\mathfrak{T}$ is not a symmetric matrix and that its indices belong to different spaces: the first labels the barycentric component and the second the Cartesian coordinate. The triangle anisotropy tensor $Q$ defined above is given by $Q=\mathfrak{T}^{T}\mathfrak{T}$. We now carry out a singular value decomposition (SVD): [^2] This decomposition was used in Ref. [36] to quantify tissue strain rates from a cell-centroid-based triangulation. However, the information contained in the “LTC phase” $\psi$ was not utilized there.
$\mathfrak{T}=R(\psi)\cdot M\cdot\begin{pmatrix}\sqrt{s_{1}}&0\\ 0&\sqrt{s_{2}}\end{pmatrix}\cdot R^{T}(\phi),$ (8)
where $R(\alpha)$ is the rotation matrix with angle $\alpha$, $M$ is a reflection matrix, the singular values $s_{1}>s_{2}>0$ by convention, and $s_{1}+s_{2}=1$ because we have normalized the tension vectors. Because of triangle symmetries, we can always choose $\psi\in[0,\pi/3]$ (see below). We now introduce the complex order parameter
$\tilde{s}\,e^{i\tilde{\psi}}\;\text{with}\;\tilde{s}=(s_{1}-s_{2})\in[0,1],\;\tilde{\psi}=3\psi\in[0,\pi]$ (9)
which provides a complete description of the triangle’s intrinsic shape. The magnitude $\tilde{s}=s_{2}-s_{1}$ measures the triangle’s anisotropy, and the “LTC phase” $\tilde{\psi}$ measures the triangle’s degree of acuteness vs obtuseness. Applied to tension triangles, the thus-defined LTC order parameter informs about the configuration of tensions at a vertex. We note that $\det M=\pm 1$ indicates the chirality of the triangle’s shape, which might be useful to detect chiral symmetry breaking on the cellular scale, e.g. in the Drosophila hindgut [37].
Geometrically, the SVD of $\mathfrak{T}$ can be understood as a sequence of transformations that map an equilateral reference triangle (with one edge parallel to the $x$-axis) to the target triangle (Fig. 4B). The reference triangle is first rotated by an angle $\psi$, then stretched along the $x$- and $y$-axis with factors $s_{1}$ and $s_{2}$, and finally rotated again by an angle $-\phi$ (the minus sign results from the convention in defining SVD). $\phi$ therefore determines the triangle’s axis of anisotropy while $\psi$ determines the orientation of the shear axis to one of the edges of the original equilateral triangle. For $\psi=0$, the shear is orthogonal to a triangle side, yielding an isosceles obtuse triangle. Increasing $\psi$ makes the resulting triangle more acute, until maximal acuteness is reached at $\psi=\pi/3$. For an illustration of how $\tilde{s},\tilde{\psi}$ relate to the internal triangle angles, see Fig. S4.
@@
Before we use the LTC parameter to quantify the tension space dynamics from simulations and experiments, we define a condition in the LTC space for when a T1 transitions will happen. This T1 threshold puts a constraint on the local tension configurations that we expect to observe. Moreover, it will allow us to quantify how tension dynamics causes active T1s by driving the local tension configurations towards the T1 threshold.
@@
where $T$ length of the shared triangle edge interface and $\ell_{0}$ fixes the length scale such that $\ell_{0}$ is the edge length of a regular hexagonal cell, corresponding to equilateral tension triangles with $T=1$. $\ell_{\mathrm{ref}}$ changes sign at $\beta+\beta^{\prime}=\pi$, which gives the “Delaunay condition” $\beta+\beta^{\prime}<\pi$. In the absence of isogonal strain, a cell neighbor exchange (corresponding to an edge flip in the triangulation) occurs at this threshold. In Fig. 4C, the gray line indicates this threshold for a pair of identical triangles (i.e. $\beta=\beta^{\prime}=\pi/2$). Notably, the threshold is at a much smaller anisotropy magnitude $\tilde{s}$ for tension cables (small $\tilde{\psi}$) than for bridges (large $\tilde{\psi}$), implying that tension cables are less efficient at driving intercalations than tension bridges.
How does the Delaunay condition generalize in the presence of isogonal strain? The length of the central interface, $\ell$, can be decomposed into two contributions
@@
where the isogonal contribution $\Delta\ell_{\text{iso}}$ accounts for isogonal modes while the (Voronoi) reference length is given by Eq. (10). Note that $\Delta\ell_{\text{iso}}$ not an edge-autonomous quantity but depends on the isogonal mode (parametrized by the isogonal function) in the four cells surrounding the interface. In practice, $\Delta\ell_{\text{iso}}$ can be estimated from the average isogonal strain tensor in a local tissue patch [17]. Now an interface collapses if the physical length reaches zero: $\ell_{\text{ref}}+\Delta\ell_{\text{iso}}=0$. This generalizes the Delaunay condition.
To find the resulting T1 threshold in the LTC space, we need to express the tension $T$ on the cell quartet’s inner interface in terms of the angles $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$. Reasoning that on average the two tension triangles in a quartet will be similar, we consider the simplified case $\alpha=\alpha^{\prime},\beta=\beta^{\prime}$. Normalizing by the average tension we find $T=3\sin(\beta)/(\sin\alpha+\sin\beta+\sin(\alpha+\beta))$. Since the problem is now reduced to the shape of one triangle, the T1 condition $\ell_{\mathrm{ref}}(\alpha,\beta)+\Delta\ell_{\mathrm{iso}}=0$ defines a threshold line in the LTC space. Figure 4E shows the T1 threshold as a function of the isogonal strain $\Delta\ell_{\mathrm{iso}}/\ell_{0}$. The critical tension anisotropy drops to zero as $\Delta\ell_{\mathrm{iso}}/\ell_{0}$ approaches $-1$, the critical isogonal strain for purely passive T1s which takes place for isotropic tension (i.e. equilateral tension triangles). Vice versa, positive isogonal strain shifts the T1 threshold to higher magnitude of tension anisotropy. In principle, the above geometric reasoning can be generalized to an arbitrary tension “kite” composed of two different tension triangles. However, the shape space of such kites is four dimensional (since there are four independent angles $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$) which precludes the intuitive visualization that the single-triangle LTC space provides.
@@
Figure 5: Dynamics of LTC order in simulations and the Drosophila germ band. A Heat maps showing the distribution of local tension configurations in a simulation with imposed isogonal strain matching the experimental observations. Data aggregated from $N=6$ simulation runs of $\sim 10^{3}$ cells each. B Driven by winner-takes-all feedback, the magnitude of tension anisotropy magnitude $\tilde{s}$ (top) and the tension bridge fraction (measured by the LTC phase $\tilde{\psi}$l; bottom) increase. The bridge fraction decreases as the tissue becomes disordered due to cell rearrangements. Solid lines show the median of the LTC distributions where the phase $\tilde{\psi}$ is weighted with the magnitude $\tilde{s}$. The width of bands showing the standard error is comparable to the line width. Shifting the median from simulations by a constant offset (see dashed orange lines) yields a quantitative match to the experimental data. C Shape dynamics of a single tension triangle driven by winner-takes-all feedback rapidly drives the tensions towards the T1 threshold with a slight bias towards tension bridges. C’ LTC distribution predicted by single-triangle simulation reproduces the LTC distribution at the onset of cell rearrangements in full tissue simulations (cf. distribution at 8 min in A). D A random Delaunay triangulation seeded from hard disks with a packing fraction $\rho=0.2$ reproduces the late-time distribution in simulations.

### v3

@@ Methods
This concludes the description of the computational model. A brief overview of the parameters and their effects is given in SI Table 1 and Fig. S9. Fig. S15 provides a flow chart of the simulation algorithm.
Figure 2: Extent of tissue flow depends on initial cell-scale order and tension anisotropy. A, B Simulation snapshots of tissue patches with free boundaries starting from an (irregular) hexagonal cell array (A) and a disordered cell array (B) generated from a random Voronoi tessellation. As small initial tension anisotropy orients convergent-extension flow. Snapshots show the top half of the tension triangulation and the bottom half of the corresponding cell array. C CE (measured by the tissue aspect ratio $a(t)$) is slower and ceases at a smaller net deformation for a disordered initial condition (cyan line) compared to the initially ordered case (purple line). Shaded bands indicate standard deviation over $N=3$ simulation runs with $N_{\mathrm{cells}}\approx 10^{3}$ cells each. D The total aspect ratio change $a_{\mathrm{final}}/a(0)$ increases as a function of cell packing order (hexagon fraction controlled by hard disk packing fraction $\rho$ used to generate the Voronoi seed points, see bottom plot) and initial tension anisotropy (controlled by a shear $s$ applied to the tension triangulation). The red star indicates the approximate state of the Drosophila germ band at the onset of germ band extension (cf. Figs. S1B, S2), corresponding to a 1.85×1.85$\times$1.85 × elongation. E Decreasing hexagon fraction (solid lines) and global tension anisotropy (dashed lines) indicate the decay of order in the cell arrays. Notably, in the initially ordered cell array, tension anisotropy transiently increases due to positive tension feedback (solid purple line). At late times both simulations converge to zero global tension anisotropy and identical coordination number statistics (inset). F Plotting the rate of tissue CE against the T1 rate provides a measure for the efficiency of T1 transitions. For the initially ordered tissue, the efficiency of T1s starts out near the theoretical optimum ($\tfrac{1}{2}\log{3}$ elongation for one T1 per cell as illustrated in the inset cartoon) but drops to zero as the tissue becomes disordered.
@@ Results
In the companion paper [Brauns.etal2024elife], we have shown that positive tension feedback can drive active T1 transitions in a regular lattice of cells with an initial anisotropy of tension. Any real tissue will exhibit some degree of irregularity. Therefore, investigating the effect of this disorder is key to understanding CE on the tissue scale. To this end, we perform simulations of freely suspended irregular cell arrays. All parameters are set to the same values as in the companion paper, where they were calibrated to fit the tension and interface length dynamics of active T1s during Drosophila gastrulation (see Fig. S6).
We generated initial tension triangulations from random hard disk packings at different packing fractions $\rho$ [Bernard.etal2009]. At low packing fraction, the hard disk process generates highly irregular triangulations (see Fig. 2B) while at sufficiently high packing fraction $\rho\gtrsim 0.72$, the disks adopt a crystalline packing such that the fraction of cells with six neighbors $p_{6}\approx 1$ (Fig. 2A). To introduce a specified initial tension anisotropy, the triangulation is sheared with magnitude $s$ (displacing vertices by $\mathbf{t}_{i}\mapsto\mathrm{diag}(\sqrt{1-s},1/\sqrt{1+s})\,\mathbf{t}_{i}$).
To quantify tension anisotropy, we define the tensor $Q=\tfrac{2}{3}\sum_{\alpha=1}^{3}\mathbf{T}_{\alpha}\otimes\mathbf{T}_{\alpha}$ for each triangle directly from the tension geometry. Averaging its deviatoric part $\tilde{Q}=Q-\frac{1}{2}\mathrm{Tr}\,Q$ over the cell array, provides a measure of global tension anisotropy $||\langle\sqrt{2}\tilde{Q}\rangle||\in[0,1]$. Starting with a slightly perturbed hexagonal cell packing and a small initial tension anisotropy, the tissue patch undergoes CE, elongating perpendicular to the initial orientation of global tension anisotropy (Fig. 2A).
@@
The tissue flow is driven by self-organized cell rearrangements (active T1 transitions) whose rate rapidly increases, reaching a maximum, and then decreases to a lower, but non-zero, value (Fig. 2F). Large-scale tissue deformation stalls after approximately 2-fold CE (as measured by the square root of the aspect ratio $a=\mathrm{width}/\mathrm{height}$, Fig. 2C) while cells continue rearranging. T1s at this stage are no longer coherently oriented and therefore do not contribute to net tissue deformation (Fig. 2F).
@@
By contrast, initializing the simulation with a low level of order in the initial cell packing, but identical tension anisotropy, results in slower convergent-extension flow and arrest of flow at a smaller amount of total tissue-scale deformation (Figs. 2B and 2C and Movie 2). Notably, tension anisotropy rapidly vanishes without the transient increase observed in the simulation starting with a more ordered cell packing (Fig. 2E). While the early dynamics depends sensitively on the initial condition, we find rapid convergence toward a common disordered steady state.
The heatmap in Fig. 2D shows the dependence of CE on the initial configuration as controlled by $\rho$ and $s$. The total extent of CE, quantified by the net change in aspect ratio $\sqrt{a_{\mathrm{final}}/a_{\mathrm{initial}}}$ increases as a function of the initial order and the magnitude of tension anisotropy. The degradation of order through cell rearrangements means that the system dynamically traverses the phase space spanned by order $p_{6}$ and anisotropy $||\langle\sqrt{2}\tilde{Q}\rangle||$ and the remaining extension is predicted by the instantaneous value of these two quantities (see SI Fig. S5).
We find that all simulations converge to a disordered state where T1s are incoherent and tissue flow stalls. This naturally explains key aspects of germ-band extension in the Drosophila embryo, in particular, the transition from the fast to the slow phase of germ-band extension [Brauns.etal2024elife], concomitant with an increase in cell-scale disorder, approaching a maximally disordered state [Irvine.Wieschaus1994, Brauns.etal2024elife] (see experimental data in Fig. S1). The self-limiting character of CE driven by positive tension feedback is robust across variations of the tension dynamics model (see SI Fig. S9, SI Sec. 2, and SI Table 1) and a similar phenomenon was observed in a recent model by Sknepnek et al. [Sknepnek.etal2023].
@@
Figure 3: Triangle shape space characterizes the local tension motifs that underlie cell rearrangements. A Local tension configurations, “cables” and “bridges”, can be quantified in terms of the tension triangle shapes. B Shape decomposition of a triangle into a sequence of three transformations acting on an equilateral reference triangle. The angle $\psi$ and the stretch factors $s_{1,2}$ determine the intrinsic shape of the triangle while the angle $\phi$ determines its orientation in space. C Intrinsic shape space of triangles parameterized by anisotropy (LTC magnitude) $\tilde{s}$ the “LTC phase” $\tilde{\psi}$ that distinguishes obtuse and acute triangles. The gray line indicates the Delaunay condition for a pair of identical triangles. Along this line, the circumcenters of the two triangles coincide, corresponding to a fourfold vertex in the Voronoi tessellation. D Circumcircle construction of the Voronoi edge length $\ell_{\mathrm{ref}}$ (purple) from a pair of adjacent (tension) triangles (red). Circumcircles are indicated by gray dashed lines. The actual physical length $\ell$ is the sum of the Voronoi reference length $\ell_{\mathrm{ref}}$, and a contribution from isogonal strain $\Delta\ell_{\mathrm{iso}}$. In the illustration, $\Delta\ell_{\mathrm{iso}}/\ell_{0}$ is negative. E T1 threshold as a function of the isogonal strain $\Delta\ell_{\mathrm{iso}}/\ell_{0}$. Positive isogonal strain shifts the threshold to higher tension anisotropy.
So far, we have focused on the role of tension anisotropy and initial topological order in the cell packing. In the Drosophila germ band, we additionally observed a more subtle form of geometric order – a particular pattern of alternating high and low tensions [Brauns.etal2024elife] – that arises dynamically before the onset of cell rearrangements. The elementary motifs of a cell-scale tension pattern are the tension triangles at individual vertices. Acute triangles correspond to tension cables (adjacent high-tension interfaces) while obtuse triangles – to which we refer as tension “bridges” – are the elementary motif of an alternating pattern of high and low tensions (Fig. 3A).
@@
To quantify the relative abundance of these motifs and compare our simulations to experimental data we define a local tension configuration (LTC) order parameter that measures how anisotropic and how acute vs obtuse a given tension triangle is. To construct this order parameter, the three tension vectors $\mathbf{T}_{\alpha},\alpha=1,2,3$ that form the tension triangle are first ordered by increasing length, i.e. $T_{1}\leq T_{2}\leq T_{3}$, and then combined into a 2$\times$2 matrix
@@
The normalization factor $\mathcal{N}$ ensures $||\mathfrak{T}||^{2}=\mathrm{Tr}[\mathfrak{T}\mathfrak{T}^{T}]=1$, fixing the arbitrary overall tension scale. $\mathfrak{T}$ is not a symmetric matrix and its indices belong to different spaces: the first labels the barycentric component and the second the Cartesian coordinate. We now carry out a singular value decomposition (SVD, geometrically illustrated in Fig. 3B): [^3] A similar decomposition was used in Ref. [Merkel.etal2017] to quantify tissue strain rates from a cell-centroid-based triangulation. However, the information contained in the “LTC phase” $\psi$ was not utilized there.
$\mathfrak{T}=R(\psi)\cdot\begin{pmatrix}\sqrt{s_{1}}&0\\ 0&\sqrt{s_{2}}\end{pmatrix}\cdot C\cdot R^{T}(\phi),$ (7)
where $R(\alpha)$ is the rotation matrix with angle $\alpha$, the singular values are ordered $s_{1}>s_{2}>0$ by convention, and $s_{1}+s_{2}=1$ because we have normalized the tension vectors. The reflection matrix $C=\mathrm{diag}(1,\pm 1)$ accounts for the chirality of the tension triangle (i.e. whether the edges go clockwise or counterclockwise if sorted by length). The angle $\phi$ represents the orientation of tension anisotropy in physical space. Indeed, the triangle anisotropy tensor $Q$ defined above is given by $Q=\mathfrak{T}^{T}\mathfrak{T}$. The intrinsic rotation angle $\psi\in[0,\pi/6]$ controls whether the subsequent shear $\mathrm{diag}\,(\sqrt{s_{1}},\sqrt{s_{2}})$ makes the triangle obtuse or acute. The intrinsic shape properties of the triangle are therefore parametrized by the magnitude of anisotropy $\tilde{s}:=(s_{1}-s_{2})\in[0,1]$ and the “LTC phase” $\tilde{\psi}=6\psi\in[0,\pi]$ mapping out a two-dimensional shape space, which we refer to as LTC space (Fig. 3C).
The SVD Eq. [7] links two spaces with different symmetries, the hexatic symmetry of the tension triangulation with the nematic symmetry of deviatoric stress in physical space. Thus, $\mathfrak{T}$ represents a hexanematic field. In contrast, the hexanematic cross-correlation defined in [Armengol-Collado.etal2024], $\mathfrak{T}$ is purely local and does not depend on the magnitude of hexatic order. Moreover, the nematic order represents the orientation of deviatoric stress, not cell elongation as in [Armengol-Collado.etal2023, Armengol-Collado.etal2024]. This distinction is important because cell elongation is controlled by the isogonal modes and can therefore decouple from stress anisotropy.
@@
The tight coupling between tension space and physical space allows us to define a condition for the occurrence of T1 transitions in LTC space. This T1 threshold will allow us to quantify how tension dynamics causes active T1s by driving the local tension configurations towards the T1 threshold. It also puts a constraint on the local tension configurations that we expect to observe.
@@
where $T$ is the length of the shared triangle edge interface and $\ell_{0}$ fixes the length scale such that $\ell_{\mathrm{ref}}=T$ for equilateral tension triangles. $\ell_{\mathrm{ref}}$ changes sign at $\beta+\beta^{\prime}=\pi$, which gives the “Delaunay condition” $\beta+\beta^{\prime}<\pi$. In the absence of isogonal strain, a cell neighbor exchange (corresponding to an edge flip in the triangulation) must occur upon crossing this threshold. In Fig. 3C, the gray line indicates this threshold for a pair of identical triangles (i.e. $\beta=\beta^{\prime}=\pi/2$). Notably, the threshold is at a much smaller anisotropy magnitude $\tilde{s}$ for tension cables (small $\tilde{\psi}$) than for bridges (large $\tilde{\psi}$), implying that tension cables are less efficient at driving intercalations than tension bridges.
How does the Delaunay condition generalize in the presence of isogonal strain? The length of the central interface, $\ell$, can be decomposed as
@@
where the isogonal contribution $\Delta\ell_{\text{iso}}$ accounts for isogonal modes while the (Voronoi) reference length is given by Eq. [8]. Note that $\Delta\ell_{\text{iso}}$ is not an edge-autonomous quantity but depends on the isogonal mode (parameterized by the isogonal function) in the four cells surrounding the interface. In practice, $\Delta\ell_{\text{iso}}$ can be estimated from the average isogonal strain tensor in a local tissue patch [Brauns.etal2024elife]. Now an interface collapses if the physical length reaches zero: $\ell_{\text{ref}}+\Delta\ell_{\text{iso}}=0$. This generalizes the Delaunay condition. Figure 3E shows the shifted T1 threshold as a function of the isogonal strain $\Delta\ell_{\mathrm{iso}}/\ell_{0}$ (see SI for a mathematical expression).